# Numerical Hopf–Lax formulae for Hamilton–Jacobi equations on unstructured geometries

### v1

@@ 6.4 Parallel implementation on GPU and complex geometries
Semi-Lagrangian schemes provide a concrete example of embarrassingly parallel algorithms, meaning that the workload for the numerical solution of our Hamilton–Jacobi equations can be naturally split among the computational resources. Indeed, referring to Algorithm 1 for evolutive problems and to the value iteration approach for stationary problems, each node of our triangular grid can be assigned to a single thread in a parallel computation, which can perform its walks along the mesh independently, using information from the previous time step or iteration. Synchronization is needed only at the end of each iteration, but since the length of each walk is similar for each node, the amount of operations is about the same for each thread, and this clearly results in a relevant speed-up of the whole computation.
Here, we implement a parallel version of the value iteration algorithm for stationary problems, written in CUDA and running on a GPU server equipped with a Nvidia H100 GPU with 94Gb Ram and 14592 CUDA cores. In the following tables, we report the results obtained for test 3 and test 4 of Section 6.2, using the same parameters and the trapezoidal quadrature rule for the approximation of the integral in the Hopf–Lax foumula, showing in particular the speed-up with respect to the serial version of the code.
$\Delta x$ $N$ CPUtime GPUtime Speed-up
0.1 394 1.33 0.07 20x
0.05 549 9.26 0.13 68x
0.025 766 66.36 0.26 256x
0.0125 1067 478.70 1.04 462x
Table 9: Value iteration method for test 3, CPU vs GPU
$\Delta x$ $N$ CPUtime GPUtime Speed-up
0.1 394 1.35 0.07 20x
0.05 549 9.38 0.13 73x
0.025 766 66.91 0.26 277x
0.0125 1067 484.76 1.02 474x
Table 10: Value iteration method for test 4, CPU vs GPU
@@
We observe that the serial version of the code does not perform well even on a relatively coarse mesh (e.g. $\Delta x=0.0125$ corresponds to a triangulation with 125189 and 249241 triangles), while the parallel version achieves an impressive speed-up. However, a key factor limiting performance is the lack of spatial locality in unstructured grids. Unlike structured meshes, where neighbouring nodes are stored contiguously in memory, an unstructured mesh requires each node to retrieve solution values and connectivity information from global memory, leading to inefficient memory access patterns. Ideally, one could improve performance by grouping spatially close nodes into thread blocks so that neighbouring nodes share cached data, reducing redundant global memory accesses. The implementation of this strategy, however, falls outside the scope of the present study.
@@ 7 Conclusions
In this paper, we introduce a numerical scheme for solving Hamilton–Jacobi equations on unstructured triangular grids using a Semi-Lagrangian approach. The key advantage of our method is that it avoids the computationally expensive interpolation step by relying solely on nodal values, making it particularly efficient for unstructured geometries.
We provide a theoretical convergence analysis within the Barles–-Souganidis framework and introduce a quadratic refinement technique to improve accuracy while maintaining computational feasibility. For stationary problems, we implemented a fast policy iteration solver, which significantly reduce computational times.
Extensive numerical tests show reliable convergence rates and robust performance even in the case of nonsmooth solutions. Additionally, we explore a GPU implementation, demonstrating substantial speed-up gains for large-scale problems.
@@ 8 Aknowledgements
S.C. is supported by the PNRR-MUR project Italian Research Center on High Performance Computing, Big Data and Quantum Computing and by the INdAM-GNCS Project code CUP E53C23001670001. R.F. and G.T. are also members of INdAM-GNCS.

### v2

@@ 6.4 Fast solvers, GPU implementation and complex geometries
Semi-Lagrangian schemes provide a concrete example of embarrassingly parallel algorithms, meaning that the workload for the numerical solution of our Hamilton–Jacobi equations can be naturally split among the computational resources. Indeed, referring to Algorithm 1 for evolutive problems and to the value iteration approach for stationary problems, each node of our triangular grid can be assigned to a single thread in a parallel computation, which can perform its walks along the mesh independently, using information from the previous time step or iteration. Synchronization is needed only at the end of each iteration, but since the length of each walk is similar for each node, the amount of operations is about the same for each thread, and this clearly results in a significant speed-up of the whole computation.
Here, we implement and compare five solvers for the stationary problem: serial and parallel versions of the value iteration, serial and parallel versions of the modified policy iteration, and serial version of the exact policy iteration. The parallel codes are written in CUDA, running on a GPU server equipped with a Nvidia H100 GPU with 94Gb Ram and 14592 CUDA cores, whereas, for a fair comparison, the serial codes have been run in single-processor mode on the same platform. Note that we have given up writing an optimized parallel code for the exact policy iteration. In fact, parallel implementations of direct solvers for linear systems are typically highly problem-dependent, and with speed-ups typically not exceeding a few units (see [11, 14, 15]). Moreover, from the point of view of a practical implementation, setting-up and managing performance of GPU libraries for sparse matrices in CUDA (e.g. cusparse) is not as simple as assembling a CUDA kernel for an iterative and explicit solver.
In Table 9, we report the results obtained for test 3 of Subsection 6.3, using the same parameters and the trapezoidal quadrature in the Hopf–Lax formula, showing in particular the speed-up with respect to the serial version of the code. Note that iterative solvers achieve an impressive speed-up, even higher in the case of Value Iteration scheme. This acceleration causes iterative solvers, which are less performant in serial implementation, to outperform the Exact Policy Iteration scheme even if the latter is implemented in CUDA.
Value Iteration Exact PI Modified PI
$\Delta x$ NN CPU GPU (x) NN CPU NN CPU GPU (x)
0.1 394 1.362 0.066 (20x) 12 0.066 12 0.149 0.032 (4.6x)
0.05 549 9.351 0.126 (74x) 15 0.276 15 0.686 0.047 (14.6x)
0.025 766 66.347 0.228 (291x) 19 1.676 19 4.264 0.077 (55x)
0.0125 1067 482.065 1.012 (476x) 24 10.506 24 28.387 0.161 (176x)
Table 9: Test 3, comparison of the computational performances associated to the three solvers.
@@
However, a key factor limiting performance is the lack of spatial locality in unstructured grids. Unlike structured meshes, where neighbouring nodes are stored contiguously in memory, an unstructured mesh requires each node to retrieve solution values and connectivity information from global memory, leading to inefficient memory access patterns. Ideally, performance could be further improved by grouping spatially close nodes into thread blocks so that neighbouring nodes share cached data, reducing redundant global memory accesses. The implementation of this strategy, however, falls once again outside the scope of the present study.
@@ 7 Conclusions
In this paper, we study the unstructured implementation of a known numerical scheme for solving Hamilton–Jacobi equations using a Semi-Lagrangian approach. The key advantage of the method is that it avoids the computationally expensive interpolation step by relying solely on nodal values; our contribution here is to improve its efficiency and accuracy on unstructured space grids.
We review the theoretical accuracy and convergence analysis within the Barles–-Souganidis framework and introduce a quadratic refinement technique to improve precision while maintaining computational feasibility. For stationary problems, we compare value iteration with fast solvers of policy iteration type, discussing computational times in both serial and GPU parallel implementations, and demonstrating, in the latter case, substantial speed-up gains for large-scale problems.
Extensive numerical tests show reliable convergence rates and robust performance even in the case of nonsmooth solutions.
@@ 8 Aknowledgements
S.C. is supported by the PNRR-MUR project Italian Research Center on High Performance Computing, Big Data and Quantum Computing and by the INdAM-GNCS Project code CUP E53C23001670001. R.F. is supported by the PRIN Project 2022238YY5, “Optimal control problems: analysis, approximation and applications”. All the Authors are members of the Gruppo Nazionale per il Calcolo Scientifico of Istituto Nazionale di Alta Matematica “Francesco Severi” (INdAM–GNCS).